### v2

# WildAvatar: Web-scale In-the-wild Video Dataset for 3D Avatar Creation

Abstract
Existing human datasets for avatar creation are typically limited to laboratory environments, wherein high-quality annotations (e.g., SMPL estimation from 3D scans or multi-view images) can be ideally provided. However, their annotating requirements are impractical for real-world images or videos, posing challenges toward real-world applications on current avatar creation methods. To this end, we propose the WildAvatar dataset, a web-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different human subjects and scenes. WildAvatar is at least 10×10$\times$10 × richer than previous datasets for 3D human avatar creation. We evaluate several state-of-the-art avatar creation methods on our dataset, highlighting the unexplored challenges in real-world applications on avatar creation. We also demonstrate the potential for generalizability of avatar creation methods, when provided with data at scale. We publicly release our data source links and annotations, to push forward 3D human avatar creation and other related fields for real-world applications.
Figure 1: Examples from our WildAvatar dataset. Unlike previous laboratory datasets for 3D avatar creation, WildAvatar curates in-the-wild web videos. With $10,000+$ human subjects and scenes, WildAvatar is at least 10×10$\times$10 × richer than the previous datasets, covering a wide range of real-world human motions and appearances. It contains high-quality annotations (i.e., SMPL, camera parameters, and human segmentation masks).
1 Introduction
Table 1: Statistics on different human datasets for avatar creation. We only consider human datasets that include human appearances. [Uncaptioned image]: multi-camera system. [Uncaptioned image]: 3D scanner. [Uncaptioned image]: depth sensor. [Uncaptioned image]: inertial measurement unit. [Uncaptioned image]: professional actor. Our WildAvatar is a large-scale and in-the-wild avatar dataset collected with our designed automatic collection pipeline.
Dataset #Sub./Sce. Type Cost Dataset #Sub./Sce. Type Cost
ZJU-Mocap [53] 9/6 Lab [Uncaptioned image]+[Uncaptioned image] RenderPeople [56] 482/-- Lab [Uncaptioned image]+[Uncaptioned image]
HuMMan [9] 339/20 Lab [Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image] THuman [84] 100/10 Lab [Uncaptioned image]+[Uncaptioned image]
DyMVHumans [82] 32/45 Lab [Uncaptioned image]+[Uncaptioned image] THuman2.0 [78] 500/-- Lab [Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]
Human3.6M [28] 11/15 Lab [Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image] THuman3.0 [63] 20/-- Lab [Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]
THuman4.0 [83] 3/3 Lab [Uncaptioned image]+[Uncaptioned image] THuman5.0 [62] 10/10 Lab [Uncaptioned image]+[Uncaptioned image]
Hi4D [76] 40/10 Lab [Uncaptioned image]+ [Uncaptioned image]+[Uncaptioned image] DNA [14] 500/1187 Lab [Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]
DynaCap [21] 4/5 Lab [Uncaptioned image]+[Uncaptioned image] MultiHuman [81] 50/-- Lab [Uncaptioned image]+[Uncaptioned image]
UltraStage [85] 100/20 Lab [Uncaptioned image]+[Uncaptioned image] Actors-HQ [29] 8/21 Lab [Uncaptioned image]+[Uncaptioned image]
HUMBI [77] 772/4 Lab [Uncaptioned image]+[Uncaptioned image] NHR [73] 4/3 Lab [Uncaptioned image]+[Uncaptioned image]
AIST++ [67] 30/10 Lab [Uncaptioned image]+[Uncaptioned image] MPII [5] 8/1 Lab [Uncaptioned image]+[Uncaptioned image]
ENeRF [42] 7/4 ITW [Uncaptioned image]+[Uncaptioned image] 3DPW [68] 7/4 ITW [Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]
FreeMan [69] 40/123 ITW [Uncaptioned image]+[Uncaptioned image] SynWild [20] 5/5 ITW --
NeuMan [31] 6/6 ITW -- TikTok [30] 340/340 ITW --
IVS-Net [15] <700/<700 ITW -- WildAvatar 10k+/10k+ ITW --
3D Human Avatar creation has extensive applications in VR/AR, film-making, metaverse, etc., attracting significant attention recently. With the advent of neural radiance fields (NeRF) [49, 66], recent research aims to recover 3D avatars from 2D observations, enabling the synthesis of novel images from arbitrary viewpoints and body poses [72, 53, 51, 26, 18, 38, 25]. For example, many works focus on reconstructing animatable human models from well-annotated and calibrated multi-view images and videos [53, 51, 52, 73], and have achieved photo-realistic results in laboratory datasets [53, 56, 4]. However, the well-designed collection systems limit its scalability to real-world scenes. To mitigate the gap between the ideal laboratory setting and real-world scenarios, a line of research proposes to reconstruct human avatars from monocular videos [19, 26, 72] or even a single image [25, 15], considering these forms of data can be easily accessible in our daily life. Still, these methods require annotations from multi-view images or even 3D scans. In addition, their performances are restricted by small-scale human data, especially for generalizable creation.
Although previous arts attempted to make existing laboratory datasets more suitable for real-world applications, there still exists a domain gap between laboratory and real-world scenarios. This is down to the fact that current human datasets are mainly collected via well-designed laboratory systems, which are expensive and time-consuming to collect. As shown in Tab. 1, existing human data collections mostly rely on accurate annotations from various advanced devices, such as well-calibrated multi-view cameras [78, 63, 48, 14], depth sensors [9, 84, 78, 28], IMUs [68], or expensive scanners [56, 76], as well as specialized actors and lightstages [21, 77, 67, 69, 42]. This ideal condition is unavailable in in-the-wild scenarios (i.e., monocular web videos) or consumer applications. To tackle this problem, recent efforts are attempting to collect in-the-wild monocular human data [30, 15, 17] but rely heavily on costly manual interventions, making it difficult to scale up. Therefore, their inadequate diversity fails to meet the requirements for dealing with in-the-wild challenges of 3D avatar creation. Given the shortage of in-the-wild human data, scaling up avatar creation from real-world scenarios is worth exploring.
To facilitate the in-the-wild avatar creation, we curate WildAvatar, a large-scale in-the-wild human avatar creation dataset extracted from YouTube, with $10,000+$ different human subjects and scenes. Specifically, we collect monocular human videos from YouTube and streamline the annotation process with off-the-shelf state-of-the-art annotation methods [55, 36, 2, 34], e.g., YOLO [55], Segment Anything [34]. Without costly sensors or lightstages, we collect over $10,000$ in-the-wild human samples. WildAvatar fills the gap in large-scale in-the-wild human data collection, offering at least 10×10$\times$10 × richer human subjects than the existing human datasets.
To investigate the performance of existing avatar creation methods in real-world scenarios, we evaluate state-of-the-art avatar creation methods [53, 51, 72, 26] on our WildAvatar dataset. Compared to laboratory datasets, existing methods encounter a significant performance drop in the reconstruction quality of the WildAvatar dataset. This highlights the gap between well-annotated laboratory datasets and in-the-wild datasets for avatar creation. We also investigate the scalability of existing generalizable human avatar creation methods [18, 38, 25] with WildAvatar, which suggests a significant improvement (up to $7\%$) in real-world scenarios. We will release the WildAvatar dataset, providing the video IDs, frame IDs, extracting scripts, and annotations obtained from our pipeline. We hope that our dataset will push forward the development of 3D human avatar creation, and other related topics, e.g., human mesh estimation (HPS) [33, 36, 41, 79, 37, 35, 27] and 3D human avatar generation [10, 23, 12, 24, 70, 46]. Overall, our main contributions are summarized as follows:
  • We present WildAvatar, a large-scale avatar dataset collected from in-the-wild videos with more than $10,000$ human subjects, at least 10×10$\times$10 × larger than previous datasets. The scale-up facilitates the creation of per-subject human avatars and paves a new avenue for generalizable human avatar reconstruction.
  • We evaluate the performance of existing state-of-the-art human avatar creation methods on the collected WildAvatar, showing the critical challenges of the in-the-wild avatar creation. Our observations could inspire the key directions in the human avatar creation.
  • We will release our data collection and annotation pipeline for WildAvatar. Further data scale-up for large-scale model training will be unlocked. The open data and code could provide important insights for future avatar creation and relevant human generation models.
2 Related Work
3D Avatar Creation Datasets. Tab. 1 presents a system overview of previous avatar creation datasets. Previous datasets mainly obtain high-quality human mask and SMPL annotations with ideal laboratory systems, e.g., well-calibrated multi-view cameras [53, 9, 82, 78, 28, 63, 83, 62, 76, 14, 21, 81, 85, 29, 77, 73, 67, 48, 42, 68], depth sensors [9, 84, 78, 28, 63, 14], IMUs [68], or expensive scanners [56, 76], as well as specialized actors and lightstages [53, 56, 9, 84, 82, 78, 28, 63, 83, 62, 76, 14, 21, 81, 85, 29, 77, 73, 67, 48, 42, 68], depth sensors [9, 84, 78, 28, 63, 14]. Unfortunately, in real-world scenarios or applications, these optimal conditions are typically hard to guarantee. Towards in-the-wild avatar creation, previous efforts also attempt to collect web human videos [30, 15]. However, they rely heavily on costly manual interventions (e.g., pre-screening or mask extraction based for  [44]) and fail to scale up. And they either show little viewpoint change [30] or not public released [15]. Therefore, there are still significant demands for a public large-scale in-the-wild dataset for avatar creation.
3D Avatar Creation Methods. Existing works on avatar creation learn coarse human surfaces with parametric mesh models [6, 47, 50, 74, 57], but these explicit meshes cannot express detailed geometry or appearance. Pifu-based methods [58, 59, 22, 40, 16] represent the human body with pixel-aligned functions and require high-quality synthetic data, but fail to generalize to real-world scenarios due to the domain gap. NeRF-based approaches [53, 51, 72, 52, 19, 39, 8, 75] learn implicit human representation from high-quality videos, and achieve photo-realistic rendering in the laboratory benchmarks. In addition, generalizable models [18, 38, 25] further simplify the data demand to a single image. Nonetheless, these methods still rely on accurate annotations, which only exist in ideal laboratory environments. Recent attempts [31, 20] decompose avatars and scenes for construction at once, demanding accurate global alignments between human bodies and backgrounds that are not available in real-world videos.
SMPL Estimation in the Wild. Parametric body models (e.g., SMPL [47], SMPLX [50]) encode coarse human surfaces with pose and shape parameters for avatar creation. For example, end-to-end methods [33, 36, 79, 41, 27] estimate the SMPL parameters efficiently from a single in-the-wild image, but they are not robust with complex scenes. Subsequent works [37, 32] refine end-to-end estimation results via fitting SMPL to 2D annotations in the loop. Recent works further consider imposing temporal consistency [2, 65, 35] to smooth SMPL estimations across input in-the-wild videos. Yet, it is far from a comprehensive annotating streamline for real-world video sources.
3 WildAvatar
Our goal is to build a large-scale in-the-wild dataset for human avatar creation. To this end, we collect abundant real-world human videos from YouTube and annotate corresponding labels. Without time-consuming and high-cost manual filtering and annotation [15], we design an efficient pipeline to filter video candidates and obtain high-quality annotations. Specifically, we download human movement web videos with our automatic pipeline, as described in Sec. 3.1. Then, as described in Sec. 3.2, we perform several processing steps (e.g., filtering and annotation) on downloaded web videos. We collect a dataset of $10k+$ real-world human videos with $200k+$ frames and annotations.
3.1 Data Collection
We follow two steps to collect human movement videos: 1) We first search for a large diverse set of video candidates on YouTube. Note that video candidates could contain some noisy videos without human movement; 2) We then download candidate videos with automatic tools [3], cut them into clips, and filter out short video clips.
Downloading Video Candidates. To cover a wide range of in-the-wild human-central activities, we start from a label pool of human motion datasets [43]. Based on these human motion labels (e.g., fishing and playing tennis), we download over $100k$ video candidates from YouTube API.
Filtering Video Clips. Some collected video candidates could not meet the high-quality human avatar creation requirement. For example, human bodies may not exist (e.g., preview, occlusion) or frequently change poses (e.g., montage) in some frames of these videos. To filter out such videos, we cut video candidates into clips with SceneDetect [1] and discard video candidates with only short clips. After filtering video candidates, we obtain $460k$ video clip candidates for further processing.
3.2 Data Processing
To obtain high-quality annotations (i.e., SMPL, camera parameters, and human segmentation mask) for collected human videos, we adopt the following four stages to annotate human video clips.
Stage I: Human Bounding Box Detection and Tracking. Precise human bounding boxes are important for monocular SMPL estimation methods [36, 33, 41, 79] and human segmentation (using Segment Anything [34]). Hence, we first obtain the bounding box of human subjects with an off-the-shelf state-of-the-art Yolo detection method [55]. Then we filter out video clip candidates with either short (e.g., less than $2$ seconds) human activity sequences or low detection or tracking confidence. Finally, we only keep the tracked subject with the largest bounding box on average (i.e., key subject) over the whole video clip sequence.
Figure 2: The four-stage data processing pipeline. We first obtain the bounding box of key subjects in videos in Stage I and extract human segmentation masks in Stage II. Then the SMPL and camera parameters are coarsely estimated in Stage III and later refined in Stage IV.
Stage II: Human Segmentation Mask Extraction. Foreground segmentation masks that distinguish human subjects from the background are important for 3D avatar creation. Previous methods [15, 72] adopt the background matting [44] methods for segmentation. These methods, however, require manual intervention to select the collected videos with at least one still background image, which is time-consuming to prepare. To efficiently segment human video clips, we adopt state-of-the-art Segment Anything (SAM) [34], which only requires bounding boxes and key points for each frame obtained in previous human detection, tracking, and 2D pose estimation steps (See Fig. 2).
Stage III: Coarse SMPL and Camera Estimation. We first estimate SMPL and camera parameters frame by frame, using a state-of-the-art single-image-based human pose and shape estimation method PARE [36]. Then, we check the pose and viewpoint shifts of the remaining video clips and only retain those with considerable viewpoint shifts and human movements. The estimated coarse SMPL and camera parameters will be used as initialization for implicit human reconstruction methods [72, 53, 51] to further refine SMPL and camera parameters in Stage IV.
Due to the depth ambiguity, monocular SMPL estimation and SAM might have low overlaps in estimated human masks. We double-check their consistency by comparing SMPL projection masks with the SAM masks. Intuitively, the SMPL mask denotes the naked body, while the SAM mask contains the clothed body. Therefore, the SMPL mask should be mostly covered by the SAM mask. To guarantee high-quality annotations, we discard video clips whose SAM-masked region has small overlaps with the SMPL-masked region.
Stage IV: Refining SMPL and Camera In-the-loop. The coarse SMPL estimated in Stage I may ignore the temporal consistency of human motions. Inspired by previous SMPL annotation pipelines [37, 2, 32], we further refine SMPL and camera parameters by smoothing these parameters across the whole video clip sequence via gradient descent. We also incorporate the estimated 2D keypoints and SAM masks into a refinement loop, providing additional supervision for precise SMPL and camera parameter estimation. Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in the Fig. A1.
3.3 Data Analysis
Our collected WildAvatar contains $10,647$ real-world human video clips with $200k+$ images. We randomly split WildAvatar into training, validation, and test splits with $7k$, $1.5k$, and $1.5k$ subjects, respectively. We show data statistics of our WildAvatar in Fig. 3 and describe the details of pose, viewpoint, and clothing distribution as follows.
Figure 3: Data Analysis: (a) word cloud of the video titles in WildAvatar, (b) histograms of annotations across video clips, here we count the bounding box and human mask region in pixels, and “Range” denotes the difference between the maximum and minimum values. (c) comparison of the body pose spaces with popular laboratory human datasets, (d) comparison of the viewpoints spaces with popular laboratory human datasets, (e) resolutions of videos in WildAvatar, and (f) comparison with the previous dataset on the abundance of clothing. We introduce the SSIOU, the inverse IOU between SMPL [47, 50] masks and segmentation masks.
Poses: Various vs. Tedious. In Fig. 3 (c), we compare the pose distribution differences among WildAvatar and previous human datasets. Specifically, we visualize the top-2 components of body poses with t-SNE [54] dimension reduction. The pose distribution comparison shows that our WildAvatar contains more diverse poses than previous laboratory datasets, as laboratory datasets only contain several designed motion sequences. This illustrates the importance of our large-scale in-the-wild WildAvatar when applying the existing models to real-world scenarios.
Viewpoints: Free vs. Fixed. We compare viewpoint distribution among previous human datasets [53, 9, 84] and WildAvatar. Specifically, we visualize the 3D rotations of observed cameras relative to human subjects. As suggested in Tab. 1, previous laboratory datasets are mainly collected in indoor lightstages, where the RGB(D) cameras’ positions are fixed for whole datasets. In real-world scenarios, however, the viewpoints of humans are arbitrary. As shown in Fig. 3 (d), the viewpoint distributions in previous datasets are sparse or unbalanced, while our WildAvatar covers a more dense viewpoint distribution.
Cloth: Diversity vs. Unitary. Previous laboratory human datasets mainly contain unitary tight clothes, while complex real-world appearances (e.g., different hairstyles) and diverse loose clothes are seldom involved. Our WildAvatar involves diverse tight and loose clothes in the real world. To quantitatively measure the clothing diversity between real-world and laboratory data, we define a metric (SSIOU) that calculates the inverse IOU between SMPL [47, 50] projected masks and human segmentation masks among different datasets in Fig. 3 (f). For humans with tight clothes, the measured SSIOU is close to $1.0$ (blue in Fig. 3 (f)), indicating their SMPL masks are similar to human segmentation masks. Correspondingly, a larger SSIOU (red in Fig. 3 (f)) reveals that human clothes are looser, with segmentation masks being larger than SMPL masks. As shown in Fig. 3 (f), more than $30\%$ subjects in WildAvatar have SSIOU values over $1.4$, suggesting that WildAvatar covers various types of clothes. This analysis validates that WildAvatar satisfies data distributions in real-world scenarios.
4 Experiments
We evaluate previous state-of-the-art methods on our WildAvatar dataset with the following commonly used metrics (details in Sec. 4.1) on 3D avatar creation. To illustrate the unclaimed challenges for avatar creation in the wild, we benchmark previous popular per-subject methods in Sec. 4.2. To further demonstrate the benefit of the large-scale WildAvatar, we compare generalizable avatar creation methods in Sec. 4.3.
4.1 Evaluation Metrics
To quantitatively evaluate the quality of novel view and novel pose synthesis, we apply three commonly used metrics: peak signal-to-noise ratio (PSNR) [60], structural similarity index (SSIM) [71], and Learned Perceptual Image Patch Similarity (LPIPS) [80]. Following previous literature [38, 25], we compute whole-image metrics for per-subject reconstruction methods, while reporting the metrics based on projected 3D human bounding box areas for generalizable methods.
4.2 Track A: Per-Subject Human Avatar Creation
Table 2: Quantitative comparisons on WildAvatar benchmark. We report the quality of novel pose synthesis of popular methods, on both the lab benchmark ZJU-Mocap and our WildAvatar.
Method Train FPS WildAvatar ZJU-Mocap
[…]
NHR [73] 5h 1.23 18.72 91.1 10.1 28.25 89.2 12.2
NeuralBody [53] 10h 1.48 16.55 85.8 12.6 30.66 96.4 4.3
Animatable NeRF [51] 10h 1.11 19.43 91.0 10.4 29.77 96.5 4.7
Animatable SDF [52] 10h 0.40 19.20 91.9 11.4 30.38 97.5 3.7
HumanNeRF [72] 10h 0.30 22.65 87.1 15.4 30.66 96.9 3.3
InstantNVR [19] 5m 2.20 23.01 91.9 9.9 31.01 97.1 3.8
GauHuman [26] 1m 189 24.02 92.1 9.1 31.34 96.5 3.1
Figure 4: Qualitative comparisons of popular per-subject methods on WildAvatar. NHP and NB, AN, AS, HN, IN, and GH denote Neural Human Rendering, NeuralBody, Animatable NeRF, Animatable SDF, HumanNeRF, InstantNVR, and GauHuman, respectively. Faces are blurred to protect privacy.
Setup. To explore the quality of WildAvatar and investigate the maximum reconstruction performance on monocular in-the-wild videos. We evaluate popular per-subject avatar creation methods on our WildAvatar dataset. We evaluate a total of 7 baselines: Neural Human Rendering (NHR[73], NeuralBody (NB[53], Animatable NeRF (AN[51], Animatable SDF (AS[52], HumanNeRF (HN[72], InstantNVR (IN[19], and GauHuman (GH[26]. Considering the calculating and time cost, it is impractical and inefficient to train all scenes in WildAvatar. Instead, we manually select $10$ representative human subjects from WildAvatar for benchmarking. For each subject, we randomly choose $10$ frames for training and the remaining $10$ frames are for testing. We follow the default model settings (learning rate, batch size, etc.) in official implementations and report average metrics of these baselines in Tab. 2.
Quantitative and Qualitative Analysis. As shown in Tab. 2, existing methods struggle to handle the more challenging in-the-wild WildAvatar, resulting in unsatisfactory performance. Compared to the popular ZJU Mocap [53] benchmark, there is a marginal drop of PSNR ($-10.12$), SSIM ($-8.04$), and LPIPS ($+6.47$) on average. We further present qualitative comparisons in Fig. 4, and outline three potential challenges towards in-the-wild avatar creation in the following.
Noisy Annotation. The quality of avatar creation significantly depends on the quality of annotations. Although we streamline our annotation pipeline with the best existing methods, it is unrealistic to annotate in-the-wild scenarios as accurately as in the lab. Specifically, existing monocular SMPL estimation methods may fail in challenging areas (e.g., hands and feet), and segmentation methods may also convey unstable artifacts (See GT in Fig. 4 (f)). It is an imperative field that investigates methods for reconstructing the human body from noisy in-the-wild data.
Loose Cloth. In-the-wild data features a wide variety of clothing, rather than dominant tight garments (e.g., T-shirts and shorts) in existing laboratory datasets. These loose clothes introduce non-rigid motions and ambiguous deformation weights, which pose significant challenges for avatar creation methods. As shown in Fig. 4 (a), (e), and (f), existing methods struggle to reconstruct loose garments and are prone to render fuzzy or floating regions. These garments are positional ambiguous to render in novel poses, as they are not considered in the existing body models [47, 50].
Uneven Light Condition. Most state-of-the-art methods [51, 72, 26, 53] reconstruct avatars only in the canonical space, and they warp the space from canonical to observation in novel pose synthesis. This fashion assumes a uniformity of light distribution of the target scenes, which is easy to set up in laboratory lightstages. However, real-world scenarios are usually illuminated in partial-occluded directions. This uneven light condition makes it more challenging to learn diffuse-aware avatars from monocular videos. Current avatar reconstruction models struggle to distinguish the reflection effects and cloth textures. This ambiguity may lead to dissonant lighting effects, e.g., the dress region in Fig. 4 (d) and legs in Fig. 4 (f).
4.3 Track B: Generalizable Human Avatar Creation
Table 3: Generalizability comparisons on challenging WildAvatar and laboratory benchmarks. We report the results of previous generalizable avatar creation methods on the cross-domain setting. NP/NV denotes Novel Pose/View. RP/THU/HM/ZJU/WA are short for RenderPeople [56] / THuman [84] / HuMMan [9] / ZJU Mocap [53] / WildAvatar (Ours) dataset. Novel Lab Test refers to the test split of laboratory datasets not included in the training.
Method Training Set WildAvatar Test Novel Lab Test
WA RP THU HM ZJU PSNR(NP) SSIM(NP) LPIPS(NP) PSNR(NV) PSNR(NP)
NHP [18] 17.39 62.6 33.2 20.05 19.83
18.12 66.3 31.4 20.49 20.24
17.35 62.2 33.4 20.84 20.81
18.07 66.9 31.2 21.20 21.27
MPS-NeRF [38] 18.21 74.2 24.0 18.46 18.76
18.79 76.8 22.2 18.48 18.77
18.20 74.3 23.8 21.15 21.43
19.11 77.2 22.1 21.19 21.48
SHERF [25] 18.39 78.1 20.2 20.67 20.10
19.43 80.5 19.3 20.96 20.49
18.31 77.9 20.4 19.07 18.99
19.50 80.7 19.2 19.63 19.73
Generalizable GS 17.76 81.1 22.8 20.31 20.03
18.43 81.5 22.3 21.73 21.65
17.73 81.2 22.7 21.47 21.78
18.39 81.4 22.4 22.46 22.02
Figure 5: Qualitative comparisons on the state-of-the-art generalizable avatar creation method [25]. “w/ WA” or “w/o WA” denotes training with or without WildAvatar, respectively. Faces are blurred to protect privacy.
Setup. We evaluate four state-of-the-art generalizable baselines on our large-scale WildAvatar, including 1) Neural Human Performer (NHP[18], 2) MPS-NeRF [38], 3) SHERF [25], and 4) Generalizable 3D Gaussian Splatting (Generalizable GS). Specifically, the Generalizable GS is adapted from GauHuman [26], with the Gaussians’ attributes decoded from the input features, following the approaches in  [11, 45]. To evaluate the overall generalizability among various scenarios, we also report results on four laboratory datasets, including 1) ZJU-Mocap [53], 2) RenderPeople [56], 3) HuMMan [9], and 4) THuman [84]. We split these datasets into training and test sets as in SHERF [25]. To avoid overfitting to a small indoor dataset, we adopt the cross-domain testing strategy, e.g.we mask out the ZJU training set when testing on the ZJU.
Quantitative and Qualitative Analysis. There are three main observations in Tab. 3. First, large-scale WildAvatar is essential for improving generalizability towards in-the-wild scenarios. Comparing odd and even rows, the WildAvatar training set brings considerable improvements in real-world scenarios. On the WildAvatar test set, it improves PSNR, SSIM, and LPIPS by 4.52%, 3.53%, and 5.06% on average, respectively. This suggests that using large-scale in-the-wild data is important for the generalizable avatar creation methods. As shown in Fig. 5, models without (w/o) training on WildAvatar, tend to predict dark-colored artifacts on the lower body. And models with (w/) training on WildAvatar can provide more realistic appearances. Second, large-scale in-the-wild data is also beneficial for laboratory benchmarks, although there are large domain gaps and the annotation of in-the-wild data is not as accurate as in the laboratory. By introducing WildAvatar dataset, we obtain an improvement of 2.57% and 2.50% PSNR on novel view and pose synthesis, respectively. Third, although 3D Gaussian Splatting (3DGS) performs well in per-subject human avatar creation, it shows slightly worse than NeRF in this generalizable setting. Our insight is that 3DGS highly relies on the point cloud initialization [26, 13, 45]. Unfortunately, due to the depth ambiguity, it is challenging to infer accurate point cloud initialization from a single image.
5 Conclusions
We curate WildAvatar, a large-scale in-the-wild human avatar creation dataset from YouTube, with 10, 000+ various human subjects and scenes. Compared with traditional avatar creation datasets, our WildAvatar consists of at least 10×10$\times$10 × more subjects or scenes. We benchmark state-of-the-art human avatar creation methods on WildAvatar, illuminate the potential of the avatar creation task with data at scale, and reveal its remaining challenges towards real-world applications. We hope this dataset will shed insights on following in-the-wild human avatar creation methods, and benefit 3D/4D human content generation works.
Pipeline Efficiency. The efficiency of data collection and annotation is also crucial to facilitate data scale-up and application. Despite the complex multiple-stage design, our data-collecting pipeline only takes less than $200$ seconds to generate full annotations for an arbitrary in-the-wild video.
Limitations. Despite the advantages of WildAvatar, we would like to list several limitations: 1) WildAvatar focuses on the 3D modeling of human bodies, lacking the 3D information of background environments. 2) To meet the needs of the avatar creation, WildAvatar contains video scenes involving dynamic human motions. Human poses in static activities (e.g., reading, fishing) are rarely included. 3) Multi-human interactions within the identical scenes are not covered.
Societal Impacts. In-the-wild avatar creations could enable potentially harmful uses. For instance, the unauthorized creation of avatars may violate portrait rights. Additionally, models trained on WildAvatar might exhibit biases in YouTube videos. Users should consider this when deploying, analyzing, and developing these models.
Appendix A Details in Data Pipeline
We provide more details on the data processing pipeline in addition to the main paper.
Stage I: Human Bounding Box Detection and Tracking. We first obtain the bounding box of human subjects with an off-the-shelf state-of-the-art Yolo detection method [55]  [^2] The configuration of the detection model can be found here.. We only keep the video clip with at least one “person” instance with its detection threshold over $0.8$. The tracking step is based on finding the largest IOU overlay of bounding boxes among frames. We discard low-resolution human subjects whose bounding box areas are lower than $64\times 64$. To ensure the richness of the dataset, we only keep one “key subject” for each video, as clips of the same video may probably share the same key subject.
Stage II: Human Segmentation Mask Extraction. To prepare key point exemplars for the Segment Anything, we first obtain the 2D keypoints $J_{2D}$ for human subjects using the popular HRNet [64]  [^3] The configuration of the 2D keypoint estimation model can be found here.. Given the 2D keypoint annotations, we are also able to discard over part-occluded subjects. Specially, we only keep the subject with the average confidence of 2D keypoints over $0.45$. For segmentation, we feed the 2D bounding box and the 2D keypoints into the sam_vit_h sub-model to extract the foreground mask.
Stage III: Coarse SMPL and Camera Estimation. We first estimate SMPL and camera parameters frame by frame, using a state-of-the-art single-image-based human pose and shape estimation method PARE [36]. To achieve better performance in complex scenes in the wild, we adapt the model pre-trained on the in-the-wild 3DPW dataset [68]. The PARE infers human body pose/shape parameters ($\theta$/$\beta$), and the global camera parameters (rotation matrix $R$ and the 3D offset $T$).
To retain the remaining video clips with considerable viewpoint shifts and human movements, we discard the clips with viewpoint angle changes lower than $\frac{\pi}{4}~{}rad$. And we automatically select $N=20$ most non-trivial frames, which keep the pose diversity to the greatest extent possible.
As mentioned in the main paper, we double-check the consistency between the SAM masks and the SMPL masks. Intuitively, the SMPL mask denotes the naked body, while the SAM mask contains the clothed body. Therefore, the SMPL mask should be mostly covered by the SAM mask (See Fig. A1 (a)  ~  (d)). We discard the subjects whose SAM masks are over 3×3$\times$3 × larger than their SMPL masks (See Fig. A1 (e)  ~  (h)). We also discard the subjects whose $10\%$ SMPL mask pixels are not covered by the SAM mask (See Fig. A1 (i)  ~  (l)).
Stage IV: Refining SMPL and Camera In-the-loop. For high-quality annotations, here we refine the coarse SMPL parameters ($\theta,\beta$) and camera parameters ($R,T$) obtained in Stage I. To achieve temporally smooth results, we regularize the differences in parameters between adjacent frames, which is given by
$\mathcal{L}^{s}_{\theta}=\sum_{i=1}^{N-1}\left\|\theta^{i}-\theta^{i+1}\right\|_{2},\mathcal{L}^{s}_{R}=\sum_{i=1}^{N-1}\left\|R^{i}-R^{i+1}\right\|_{2},\mathcal{L}^{s}_{T}=\sum_{i=1}^{N-1}\left\|T^{i}-T^{i+1}\right\|_{2},$ (1)
$\mathcal{L}^{s}_{2D}=\sum_{i=1}^{N-1}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i});R^{i},T^{i})-\Pi(J_{3D}(\theta^{i+1},\beta^{i+1});R^{i+1},T^{i+1})\right\|_{2},$
where $J_{3D}(\theta,\beta)$ infers the 3D keypoints of the human body, and the $\Pi$ denotes the 2D projection, and $i$ denotes the $i_{th}$ frame of the input video. Notice that the body shapes ($\beta$) are treated as constants across the input video.
Figure A1: SMPL and SAM consistency. The green/red borders denote good/bad outputs, respectively.
In addition, we align the human body parameters to the 2D keypoints and SAM masks, which are given by
$\mathcal{L}_{2D}=\sum_{i=1}^{N}\left\|\Pi(J_{3D}(\theta^{i},\beta^{i}))-J^{i}_{2D}\right\|_{2},\mathcal{L}_{\mathrm{mask}}=\sum_{i=1}^{N}d_{CD}(\Pi(V_{3D}(\theta^{i},\beta^{i})),M^{i}),$ (2)
where $d_{CD}$ denotes the chamfer distance between point clouds, and the $V_{3D}(\theta,\beta)$ infers the 3D vertex of the human body. The $M^{i}$ denotes the foreground d pixels cloud of the $i$ frame. We also regularize $\theta$ to avoid out-of-domain poses [7], using the Gaussian Mixture Model (GMM) prior [50]
$\mathcal{L}_{\mathrm{prior}}=\sum_{i=1}^{N}\left\|GMM(\theta^{i})\right\|_{2}.$ (3)
We adopt the loss functions as mentioned above for supervision:
$\mathcal{L}=\lambda^{s}_{\theta}\mathcal{L}^{s}_{\theta}+\lambda^{s}_{R}\mathcal{L}^{s}_{R}+\lambda^{s}_{T}\mathcal{L}^{s}_{T}+\lambda^{s}_{2D}\mathcal{L}^{s}_{2D}$ (4)
$+\lambda_{2D}\mathcal{L}_{2D}+\lambda_{\mathrm{mask}}\mathcal{L}_{\mathrm{mask}}+\lambda_{\mathrm{prior}}\mathcal{L}_{\mathrm{prior}}.$
We empirically set the loss weights as $\lambda^{s}_{\theta}=100$, $\lambda^{s}_{R}=1000$, $\lambda^{s}_{T}=50$, $\lambda^{s}_{2D}=100$, $\lambda_{2D}=100$, $\lambda_{\mathrm{mask}}=100$ and $\lambda_{\mathrm{prior}}=0.1$. We adopt the LBFGS [61] optimizer with the learning rate $lr=1.0$. We optimize for $20$ steps on a single Geforce GTX 1080Ti GPU.
Qualitative comparisons between coarse and refined SMPL and camera parameters can be found in Fig. A2.
Figure A2: Comparison of the coarse and refined SMPL parameters.
Appendix B Additional Links
We release the WildAvatar dataset with a permanent doi 10.5281/zenodo.11526806. We also provide the Croissant metadata. The authors bear all responsibility in case of violation of rights, and confirm that this dataset is open-sourced under the CC BY 4.0 International license.